1. 12m-4>20
2. x-2<(x+1/3)>0
3. 1/x-1<0<1/x+1
4. 0 5. l-4(3-2xl>3<<<

Answer by lenny460(1073)

1. 12m-4>20
2. x-2<(x+1/3)>0
3. 1/x-1<0<1/x+1
4. 0 5. l-4(3-2xl>3<<<
1. 12m - 4 > 20
Add 4 to each side of the inequality.
12m - 4 + 4 > 20 + 4
Combine like terms:
-4 + 4 = 0
20 + 4 = 24
Therefore:
12m = 24
Divide each side by 12
12m/12 = 24/12
m = 2




Lennox Obuong
Algebra Student
Email: obuong3@aol.com